A blind person after walking each and steps in One Direction each of length it is and turns randomly to the left to the right by 90 degree after walking a total of 40 steps the maximum possible displacement of the person from his starting position could be

Grade:12th pass

1 Answers

Arun
6 years ago
He took 10 steps of 80cm that is = 10×80 cm
=800cm = 8m
after turning right or left he took 40 steps. that is= 40×80cm=3200cm=32m
His displacement=✓8²+32²=✓64+1024=✓1088
=32.9848...m, or 32m approx.
therefore, correct option is option 2. that is 32m.